# Foundation Models for Remote Sensing and Earth Observation: A Survey

### v2

@@ Abstract
Remote Sensing (RS) is a crucial technology for observing, monitoring, and interpreting our planet, with broad applications across geoscience, economics, humanitarian fields, etc. While artificial intelligence (AI), particularly deep learning, has achieved significant advances in RS, unique challenges persist in developing more intelligent RS systems, including the complexity of Earth’s environments, diverse sensor modalities, distinctive feature patterns, varying spatial and spectral resolutions, and temporal dynamics. Meanwhile, recent breakthroughs in large Foundation Models (FMs) have expanded AI’s potential across many domains due to their exceptional generalizability and zero-shot transfer capabilities. However, their success has largely been confined to natural data like images and video, with degraded performance and even failures for RS data of various non-optical modalities. This has inspired growing interest in developing Remote Sensing Foundation Models (RSFMs) to address the complex demands of Earth Observation (EO) tasks, spanning the surface, atmosphere, and oceans. This survey systematically reviews the emerging field of RSFMs. It begins with an outline of their motivation and background, followed by an introduction of their foundational concepts. It then categorizes and reviews existing RSFM studies including their datasets and technical contributions across Visual Foundation Models (VFMs), Visual-Language Models (VLMs), Large Language Models (LLMs), and beyond. In addition, we benchmark these models against publicly available datasets, discuss existing challenges, and propose future research directions in this rapidly evolving field. A project associated with this survey has been built at https://github.com/xiaoaoran/awesome-RSFMs.
@@
Foundation model, remote sensing, geoscience, multimodal, visual recognition, vision-language model, large language model, earth observation, artificial intelligence.1 Introduction
@@
The major contributions of this work are threefold: First, it provides a thorough and systematic review of recent advancements in RSFMs. To the best of our knowledge, this is the first survey that spans different types of FMs in this rapidly evolving field. Second, it benchmarks and offers an in-depth analysis of RSFMs applied across various sensor modalities and tasks. Third, it identifies several research challenges and proposes potential research directions in the domain of RSFMs.
The structure of this survey is as follows: In Section 2, we provide background knowledge on RSFMs, including learning paradigms, common RS sensor modalities, and related surveys. Section 3 delves into the foundations of RSFMs, covering deep network architectures and typical RS interpretation tasks. Sections 4, 5, and 6 offer a systematic review of methods for VFMs in RS, VLMs in RS, and other types of RSFMs. In Section 7, we summarize and compare the performance of existing methods across multiple benchmark datasets. Finally, Section 8 outlines several promising future directions for RSFMs.
2 Background
2.1 RS Learning Paradigms
@@
2.2 Common RS Sensor Modalities
@@
2.3 Relevant Surveys
To the best of our knowledge, this is the first survey to systematically review FMs across various RS fields, including VFMs, VLMs, LLMs, generative FMs, and beyond. While numerous surveys in natural domains [1, 5] cover diverse scopes and applications, they do not specifically address geoscience and RS modalities. Additionally, several RS-specific surveys have explored topics such as foundations for earth and climate RSFMs [6], self-supervised learning [7, 8], and vision-language modeling [9]. Our survey aims to bridge these gaps by providing a comprehensive and up-to-date overview of various RSFM types and RS modalities, highlighting the latest advancements in the field.
@@ 3 Foundations of RSFMs
In Sections 4, 5 and 6, we will provide a detailed review of the pretraining methods used for RSFMs, along with their corresponding pre-training datasets. In this section, we will first introduce foundational deep architectures in subsection 3.1 and then discuss the most common downstream RS tasks in subsection 3.2.
3.1 Deep Network Architectures
@@
3.2 Typical RS Interpretation Tasks
@@
4 Visual Foundation Models for RS
4.1 VFM Pre-training in RS
4.1.1 Pre-training Datasets
@@
Table I provides an overview of the pre-training datasets utilized in recent RSFMs. A clear trend towards larger datasets is evident, with an increasing availability of training data and a broader inclusion of diverse RS modalities. It is promising to see these datasets being collected from various sources and featuring different Ground Sample Distances (GSD), creating more comprehensive databases for RS research. On the other hand, the high cost of annotation poses practical limitations on the benefits of pre-training, which has driven a shift toward Self-Supervised Learning (SSL) where pre-training datasets are unlabeled and easier to construct. Despite these advancements, RS pre-training datasets still fall short in scale—both in terms of size and modality diversity—when compared to those used in general foundation models.
Next, we review VFM pre-training methods in RS domains, including both supervised and self-supervised approaches. Table II summarizes representative methods.
4.1.2 Supervised Pre-training
@@
4.1.3 Self-Supervised Pre-training
@@
where is the query embedding, are the key embeddings, with representing the positive key for and the remaining are treated as negative keys. The hyperparameter controls the smoothness or separation of the learned representations.
While contrastive pre-training [39, 36, 40, 41, 42] has made significant contributions in computer vision, it has also inspired numerous efforts in RS, particularly within the optical RGB modality [43]. For example, GASSL [23] enhanced the MoCo-v2 framework [40] by integrating geo-location prediction as an additional pretext task. CACo [25] utilized contrastive learning to detect both short-term and long-term changes, leveraging the spatiotemporal structure of temporal RS image sequences. MATTER [24] employed multi-temporal, spatially aligned RS imagery over unchanged regions to develop representations invariant to illumination and viewing angles, specifically for materials and textures. CSP [26] introduced contrastive spatial pre-training for geo-tagged images, enriching representation learning through geo-location information. Furthermore, Huang et al. [44] combined self-supervised contrastive learning on unlabeled RS images with supervised learning on labeled natural images, demonstrating that integrating generic knowledge can significantly enhance pre-training for RS imagery.
Several studies have expanded contrastive learning to include modalities beyond RGB. For instance, SeCo [12] applied contrastive learning to MSI, enabling the learning of time- and position-invariant representations over seasonal changes. Li et al. [45] incorporated global land cover products and the geographical locations of RS images to provide additional supervision in their contrastive learning framework. Additionally, Li et al. [46] implemented both global and local contrastive learning techniques on RGB and Near-Infrared (NIR) images. SkySense [27] introduced a contrastive learning framework that integrates multiple modalities and spatial granularities, leveraging geo-context prototype learning to facilitate cross-modal processing among RGB, MSI, and SAR images.
@@
MIM has been widely applied in the RS domain; however, the original MAE framework [37], which is designed for static natural RGB images, often struggles with the unique characteristics of RS imagery. To address these challenges, several RS-specific adaptations have been proposed. For instance, to counter common small objects in RS images that random masking can overlook, Sun et al. [47] introduced a patch-incomplete masking strategy. SatMAE [13] focused on MAE pre-training for temporal MSI, while SatMAE++ [31] proposed multi-scale reconstructions for both RGB and MSI, tackling the scales variability in RS images.
To handle the varying sizes and arbitrary orientations of objects in RS imagery, Wang et al. [48] developed a transformer with rotated, varied-size window attention, paired with MAE pre-training. Alternatively, MA3E [32] pre-trains models using angle-aware MIM to learn angle-invariant representations. Some approaches also explore reconstruction in the frequency domain [34, 49]. For instance, Scale-MAE [34] combines low/high frequency reconstruction with ground sample distance positional encoding.
Beyond spatial dimensions, researchers have explored masking in the spectral dimension for MSI [50], HSI [51], and multimodal data [52]. Additionally, efforts have been made to develop efficient MAE frameworks tailored to the RS domain: Mendieta et al. [15] investigated continual pre-training using ImageNet models as an auxiliary distillation objective to enhance geospatial FMs while reducing downstream computational costs; Wang et al. [53] introduced an efficient MAE that encodes and reconstructs only a subset of semantically rich patch tokens.
@@ 4.2.2 SAM-based studies for RS
Furthermore, SAM’s design for single-modal RGB images limits its application to the multimodal data often used in RS, which typically incorporates data from multiple sensor types. MM-SAM [66] expands SAM’s capabilities to support cross-modal and multimodal processing, enabling enhanced segmentation across a variety of sensor suites commonly used in RS, such as RGB plus HSI, RGB plus LiDAR, etc.
4.3 Summary and Discussion
@@
Additionally, the introduction of SAM has spurred numerous adaptations tailored to RS data, extending its utility across different RS-specific scenarios, semantic recognition tasks, cross-modal transfer, and multimodal processing. While there are a few studies exploring other VFMs, such as depth anything model [67] for tasks like canopy height estimation [68] and building extraction [69], progress in these areas has been gradual. Significant advancements are still anticipated to unlock the full potential of VFMs in RS.
5 Vision-Language Modelsfor RS
@@
5.1 Preliminaries of VLMs in Natural Domains
@@
[…]
5.2 Vision-language datasets for RS
Building on the advancements of VLMs in general domains, the RS field is actively developing RS-specific vision-language datasets as a foundation for technical progress. It’s noted that vision-language tasks in RS have a long history, even prior the era of FMs. To offer a comprehensive overview of these efforts, we summarize the existing vision-language datasets for RS in Table III.
@@ 5.2 Vision-language datasets for RS
5.3 VLMs with Contrastive Objectives
Similar to the contrastive learning in VFMs discussed in Section 4.1.3, the contrastive pre-training in VLMs [70] also employs the InfoNCE loss [38] as the training objective. A typical example of FM is CLIP [70]. The primary difference is that, in CLIP, the contrastive samples are image-text pairs, and the objective becomes a symmetric image-text InfoNCE loss, defined as:
@@ 5.3 VLMs with Contrastive Objectives
5.3.1 Visual-Text Models
@@
5.3.2 Visual-Location Models
@@
5.4 VLMs with Generative Objectives
@@
5.5 Summary and Discussion
@@
6 Other RSFMs
@@
6.1 Large Language Models for RS
@@
6.2 Generative Models for RS
While most previous studies focus on discriminative downstream tasks, generative foundation models (FMs), such as diffusion models [133, 134, 135, 136], represent a crucial category in various image generation tasks. Trained on large-scale image-text datasets, these models can generate high-resolution images from user prompts (e.g., texts) and are applied to inverse problems like inpainting, colorization, deblurring, and video generation.

### v3

@@ Abstract
Remote Sensing (RS) is a crucial technology for observing, monitoring, and interpreting our planet, with broad applications across geoscience, economics, humanitarian fields, etc. While artificial intelligence (AI), particularly deep learning, has achieved significant advances in RS, unique challenges persist in developing more intelligent RS systems, including the complexity of Earth’s environments, diverse sensor modalities, distinctive feature patterns, varying spatial and spectral resolutions, and temporal dynamics. Meanwhile, recent breakthroughs in large Foundation Models (FMs) have expanded AI’s potential across many domains due to their exceptional generalizability and zero-shot transfer capabilities. However, their success has largely been confined to natural data like images and video, with degraded performance and even failures for RS data of various non-optical modalities. This has inspired growing interest in developing Remote Sensing Foundation Models (RSFMs) to address the complex demands of Earth Observation (EO) tasks, spanning the surface, atmosphere, and oceans. This survey systematically reviews the emerging field of RSFMs. It begins with an outline of their motivation and background, followed by an introduction of their foundational concepts. It then categorizes and reviews existing RSFM studies including their datasets and technical contributions across Visual Foundation Models (VFMs), Visual-Language Models (VLMs), Large Language Models (LLMs), and beyond. In addition, we benchmark these models against publicly available datasets, discuss existing challenges, and propose future research directions in this rapidly evolving field.
@@
Foundation model, remote sensing, geoscience, multimodal, visual recognition, vision-language model, large language model, earth observation, artificial intelligence.I Introduction
@@
The major contributions of this work are threefold: First, it provides a thorough and systematic review of the latest advancements in RSFMs. Second, it benchmarks and offers an in-depth analysis of RSFMs applied across various sensor modalities and tasks. Third, it identifies several research challenges and proposes potential research directions in the domain of RSFMs.
The structure of this survey is as follows: In Section II, we provide background knowledge on RSFMs, including learning paradigms, common RS sensor modalities, and related surveys. Section III delves into the foundations of RSFMs, covering deep network architectures and typical RS interpretation tasks. Sections IV, V, and VI offer a systematic review of methods for VFMs in RS, VLMs in RS, and other types of RSFMs. In Section VII, we summarize and compare the performance of existing methods across multiple benchmark datasets. Finally, Section VIII outlines several promising future directions for RSFMs.
II Background
II-A RS Learning Paradigms
@@
II-B Common RS Sensor Modalities
@@
| Pub. | Year | Scopes & Contributions |
| [5] | 2022 | Developments in SSL for CV in the context of RS. |
| [6] | 2023 | Developments of SSL for various RS applications. |
| [7] | 2024 | Foundations for earth and climate RSFMs. |
| [8] | 2024 | Progress and trends for VLMs in RS. |
| [9] | 2024 | Generative AI and LLMs for RS. |
| [10] | 2024 | Broad overview of FMs in geoscience. |
| Ours | 2025 | RSFMs in dataset development, model taxonomy, |
| technical advances, and benchmarking. |
II-C Relevant Surveys
This survey systematically review and benchmark FMs across various RS fields, including VFMs, VLMs, LLMs, generative FMs, and beyond. While numerous surveys in natural domains [1, 11] cover diverse scopes and applications (see Table I), they do not specifically address geoscience and RS modalities. Additionally, there are several surveys regarding in geoscience and RS but they focus on specific topics such as foundations for earth and climate RSFMs [7], self-supervised learning [5, 6], and vision-language modeling [12]. [10] present present a broad yet high-level review of FMs in geoscience, primarily emphasizing general-domain milestone models rather than geoscience-specific advancements. The review lacks in-depth analysis of dataset resources and technical progress in RS and geoscience, limiting its effectiveness in capturing the latest field developments. Moreover, these reviews predominantly rely on qualitative analyses, omitting the quantitative benchmarking necessary for a more comprehensive evaluation. Our survey aims to bridge these gaps by providing a comprehensive and up-to-date overview of various RSFM types and RS modalities, highlighting the latest advancements in the field.
@@ III Foundations of RSFMs
In Sections IV, V and VI, we will provide a detailed review of the pretraining methods used for RSFMs, along with their corresponding pre-training datasets. In this section, we will first introduce foundational deep architectures in subsection III-A and then discuss the most common downstream RS tasks in subsection III-B.
III-A Deep Network Architectures
@@
III-B Typical RS Interpretation Tasks
@@
IV Visual Foundation Models for RS
@@
| OmniSat [37] [code] | SSL-C | ECCV2024 | RGB, SAR | Cross-modal contrast among temporal and high-resolution RGB images and SAR images. |
@@
| HyperSIGMA [22] [code] | SSL-M | TPAMI2025 | HSI | A FM for HSI with MIM pre-training on spatial and spectral dimensions. |
| HyperFree[23] [code] | Sup. | CVPR2025 | HSI | A channel-adaptive, tuning-free FM for HSI enabling promotable and semantic-aware segmentation. |
IV-A VFM Pre-training in RS
IV-A1 Pre-training Datasets
Table II provides an overview of the pre-training datasets utilized in recent RSFMs. A clear trend towards larger datasets is evident, with an increasing availability of training data and a broader inclusion of diverse RS modalities. It is promising to see these datasets being collected from various sources and featuring different Ground Sample Distances (GSD), creating more comprehensive databases for RS research. On the other hand, the high cost of annotation poses practical limitations on the benefits of pre-training, which has driven a shift toward Self-Supervised Learning (SSL)where pre-training datasets are unlabeled and easier to construct. Despite these advancements, RS pre-training datasets still fall short in scale—both in terms of size and modality diversity—when compared to those used in general foundation models.
Next, we review VFM pre-training methods in RS domains, including both supervised and self-supervised approaches. Table III summarizes representative methods.
IV-A2 Supervised Pre-training
@@
IV-A3 Self-Supervised Pre-training
@@
where is the query embedding, are the key embeddings, with representing the positive key for and the remaining are treated as negative keys. The hyperparameter controls the smoothness or separation of the learned representations. While contrastive pre-training [45, 42, 46, 47, 48] has made significant contributions in computer vision, it has also inspired numerous recent efforts in RS that can be grouped by sensor modalities:
Optical RGB: Majority of studies focus particularly on the optical RGB modality [49] For example, GASSL [27] enhanced the MoCo-v2 framework [46] by integrating geo-location prediction as an additional pretext task. CACo [29] utilized contrastive learning to detect both short-term and long-term changes, leveraging the spatiotemporal structure of temporal RS image sequences. MATTER [28] employed multi-temporal, spatially aligned RS imagery over unchanged regions to develop representations invariant to illumination and viewing angles, specifically for materials and textures. CSP [30] introduced contrastive spatial pre-training for geo-tagged images, enriching representation learning through geo-location information. Furthermore, Huang et al. [50] combined self-supervised contrastive learning on unlabeled RS images with supervised learning on labeled natural images, demonstrating that integrating generic knowledge can significantly enhance pre-training for RS imagery.
Non-optical RGB: Several studies have expanded contrastive learning to include modalities beyond optical RGB. For instance, SeCo [15] applied contrastive learning to MSI, enabling the learning of time- and position-invariant representations over seasonal changes. Li et al. [51] incorporated global land cover products and the geographical locations of RS images to provide additional supervision in their contrastive learning framework. Additionally, Li et al. [52] implemented both global and local contrastive learning techniques on RGB and NIR images. SkySense [31] introduced a contrastive learning framework that integrates multiple modalities and spatial granularities, leveraging geo-context prototype learning to facilitate cross-modal processing among RGB, MSI, and SAR images.
@@
MIM has been widely applied in the RS domain; however, the original MAE framework [43], which is designed for static natural RGB images, often struggles with the unique characteristics of RS imagery. To this end, a series of studies have been proposed to address these RS-specific challenges based on the MIM architectures:
Varying temporal resolutions: SatMAE [16] focused on MAE pre-training with temporal embedding along with independently masking image patches across time.
Varying spatial resolutions: SatMAE++ [35] proposed multi-scale reconstructions for both RGB and MSI for addressing spatial resolution variability in RS images. Some approaches also explore reconstruction in the frequency domain [39, 53]. For instance, Scale-MAE [39] combines low/high frequency reconstruction with ground sample distance positional encoding.
Varying object sizes and orientations: Sun et al. [54] introduced a patch-incomplete masking strategy to tackle small object sizes common in RS images, which random masking can overlook. Wang et al. [55] developed a transformer with rotated, varied-size window attention, paired with MAE pre-training. Alternatively, MA3E [36] pre-trains models using angle-aware MIM to learn angle-invariant representations.
Additionally, efforts have been made to develop efficient MAE frameworks tailored to the RS domain: Mendieta et al. [18] investigated continual pre-training using ImageNet models as an auxiliary distillation objective to enhance geospatial FMs while reducing downstream computational costs; Wang et al. [58] introduced an efficient MAE that encodes and reconstructs only a subset of semantically rich patch tokens.
@@ IV-B2 SAM-based studies for RS
IV-C Summary and Discussion
@@
Additionally, the introduction of SAM has spurred numerous adaptations tailored to RS data, extending its utility across different RS-specific scenarios, semantic recognition tasks, cross-modal transfer, and multimodal processing. While there are a few studies exploring other VFMs, such as depth anything model [71] for tasks like canopy height estimation [72] and building extraction [73], progress in these areas has been gradual.
It is noted that while many of reviewed frameworks have been trained and tested in multiple sensor modalities individually, only a few RSFMs have been developed to accept multiple sensor modalities simultaneously as complementary input. For instance, msGFM [34] demonstrates that multimodal pre-training can enhance downstream performance in multimodal inference tasks following fine-tuning. Another example is TerraMind [74], which incorporates diverse modalities for paired data, including optical RGB, SAR, NDVI, LULC, DSM, text captions, and geolocation coordinates. Leveraging a globally available public dataset, TerraMind conducts large-scale multimodal MIM pre-training and achieves promising results in cross-modal synthetic data generation, as well as in zero-shot and few-shot scenarios. MM-SAM [75] expands SAM’s RGB processing capabilities to support cross-modal and multimodal processing, enabling enhanced segmentation across a variety of sensor suites commonly used in RS, such as RGB plus HSI, RGB plus LiDAR, etc. Nevertheless, the performance gains reported in these studies over single-modality fine-tuning remain relatively modest. This suggests that, while multimodal learning is a promising direction, the effective integration of heterogeneous sensor modalities continues to pose significant challenges, largely due to the intrinsic differences in spatial resolution, spectral characteristics, and physical sensing principles across modalities.
Importantly, VFMs have demonstrated new capabilities that surpass traditional RS approaches. Examples include any-to-any generative frameworks across multiple RS sensor modalities and products [74] via multimodal MIM, accelerated semantic mask annotation pipelines based on SAM [60], and achieving high segmentation performance with minimal fine-tuning data [70]. These advancements suggest that VFMs are beginning to accelerate the evolution of RS interpretation tasks. Although current RSFM studies have not yet fundamentally reshaped a specific RS application that traditionally addressed by conventional approaches, comparable to the paradigm shifts brought by GPT-4 or SAM in the general domain, the future holds substantial promise for unlocking the full potential of VFMs in RS.
V Vision-Language Models for RS
@@
V-A Preliminaries of VLMs in Natural Domains
[…]
@@ V-A Preliminaries of VLMs in Natural Domains
V-B Vision-language datasets for RS
Building on the advancements of VLMs in general domains, the RS field is actively developing RS-specific vision-language datasets as a foundation for technical progress. It’s noted that vision-language tasks in RS have a long history, even prior the era of FMs. To offer a comprehensive overview of these efforts, we summarize the existing vision-language datasets for RS in Table IV.
@@ V-B Vision-language datasets for RS
V-C VLMs with Contrastive Objectives
Similar to the contrastive learning in VFMs discussed in Section IV-A3, the contrastive pre-training in VLMs [76] also employs the InfoNCE loss [44] as the training objective. A typical example of FM is CLIP [76]. The primary difference is that, in CLIP, the contrastive samples are image-text pairs, and the objective becomes a symmetric image-text InfoNCE loss, defined as:
@@ V-C VLMs with Contrastive Objectives
V-C1 Visual-Text Models
@@
V-C2 Visual-Location Models
@@
V-D VLMs with Generative Objectives
@@
While previous efforts focus on single static image, TEOChat [107] engages in conversations about temporal sequences of earth observation data, including building change and damage assessment, semantic change detection, and temporal scene classification. EarthDial[136] is a recently proposed conversational assistant for RS, supporting multiple EO tasks including classification, detection, captioning, VQA, visual reasoning, and grounding, across multi-spectral, multi-temporal, and multi-resolution imagery. It accommodates various modalities, including RGB, SAR, and multispectral bands (e.g., NIR and infrared), and enables bi-temporal and multi-temporal analysis for change detection.
Prior studies often integrate RS imagery with varying GSD s to accommodate differences in spatial resolution; however, their tuning data is typically limited to low- and mid-resolution images. A distinct challenge in RS lies in interpreting very high-resolution imagery, where image sizes are large and objects of interest are often extremely small. To evaluate the capabilities of existing FMs in this context, Wang et al. [137] introduced XLRS-Bench, a benchmark constructed from ultra-high-resolution RS imagery, which reveals significant performance limitations of current models under such conditions.
V-E Summary and Discussion
@@
While most current VLMs in RS are still tailored to specific applications, they demonstrate early signs of enabling more flexible, scalable, and user-friendly RS solutions. These models offer a novel paradigm for RS interpretation, potentially surpassing traditional, task-specific approaches by fostering more interactive and versatile use of geospatial data.
VI Other Foundations for RSFMs
@@
VI-A Large Language Models for RS
@@
VI-B Generative Models and Data Synthesis for RS
While most previous studies on RSFMs have concentrated on discriminative downstream tasks, generative modeling represents a crucial and complementary direction, particularly for data synthesis. The development of RSFMs requires vast and diverse datasets for effective pretraining; however, RS data is often scarce, incomplete, heterogeneous in modality, and costly to collect or annotate. Synthetic data generation offers a promising solution by expanding both the diversity and quantity of available data, often providing automatic annotations and thereby accelerating the creation of large-scale, versatile RS models.
Generative foundation models, especially diffusion models [143, 144, 145, 146], have achieved remarkable success in natural image generation, typically leveraging auxiliary information such as text prompts or metadata for controlled generation. Nevertheless, RS data possesses unique characteristics, such as varing spectral bands, irregular temporal sampling, and complex spatio-temporal structures, that present additional challenges beyond those encountered in natural image domains. These differences necessitate the development of specialized generative approaches tailored to RS data properties. Despite these challenges, generative models hold significant potential for advancing RS applications, including super-resolution, cloud removal, temporal inpainting, and cross-modal data generation, underscoring the urgent need for RS-specific generative foundation models.
Recent studies have started to explore this direction. Khanna et al. [147] introduced DiffusionSat, the first generative FM specifically designed for satellite imagery, inspired by the Stable Diffusion architecture [146]. DiffusionSat conditions image generation not only on textual descriptions but also on satellite-specific metadata such as latitude, longitude, timestamps, and GSD. This metadata-driven conditioning allows the model to be flexibly adapted for diverse generation tasks, including single-image synthesis, super-resolution, inpainting, and temporal generation from publicly available satellite datasets. In parallel, Zheng et al. [148] proposed a generative probabilistic change detection model based on diffusion processes, showcasing the utility of generative models for tackling core EO problems.
In addition to diffusion-based approaches, multimodal masked modeling has emerged as another pathway for synthetic data generation. For example, TerraMesh [149] constructs a globally diverse multimodal dataset by combining optical imagery, SAR, DEM, and land-cover data, using masked reconstruction objectives to generate synthetic multimodal outputs. This enables the creation of rich, cross-modal training samples that are otherwise difficult and expensive to acquire at large scales.
Another promising direction is physics-based simulation for RS data synthesis [150, 151]. Pipelines such as SynRS3D [151] leverage 3D graphics engines like Blender, enhanced with text descriptions generated by GPT-4 and texture realism from Stable Diffusion, to synthesize large-scale, high-resolution optical datasets. SynRS3D features diverse land cover types, precise height information, and annotated building changes, offering a scalable and flexible alternative for creating high-quality RS datasets tailored for training RSFMs.
Collectively, these generative modeling efforts signal a pivotal expansion of RSFM capabilities beyond traditional discriminative tasks, highlighting the importance of synthetic data pipelines in enabling robust, versatile, and scalable remote sensing foundation models.
VI-C Computational Efficiency and Scalability
FMs are highly computation-intensive, and RS applications further exacerbate this due to high-resolution, multimodal, and temporally rich data, along with demands for real-time processing and edge deployment. These factors pose substantial challenges to the scalability and efficiency of RSFMs.
A commonly adopted strategy to mitigate training data and computational requirements is the use of PEFT methods (e.g., LoRA and adapter modules) as well as knowledge distillation. As reviewed in previous sections, most of RSFM studies have leveraged these techniques to reduce overhead while maintaining adaptability to RS-specific tasks.
Beyond fine-tuning, another promising direction lies in the design of lightweight backbone architectures tailored for RS edge-device deployment. For example, Wang et al. [152] proposed a streamlined architecture that enhances spatial context modeling in RS imagery by dynamically adjusting the spatial receptive field to accommodate object-scale variations commonly observed in RS data. Similarly, Li et al. [153] introduced a dual-branch CNN–Transformer hybrid architecture optimized for on-orbit inference, striking a balance between representational capacity and computational efficiency.
In addition, model compression remains a critical yet relatively underexplored area in RSFMs. Its primary goal is to reduce model size, inference latency, and memory footprint while preserving task performance, which is an essential step for enabling deployment in resource-constrained environments. An early effort in this direction was introduced by Shinde et al. [154].
Given the multimodal heterogeneity and spatiotemporal complexity of RS data, future efforts should target specialized strategies, including multimodal-aware compression, adaptive inference, and dynamic resource management, to enable practical and scalable RSFM deployment.
VI-D Weather Forecasting Foundation Models
VI-E Responsible RSFMs
As RSFMs continue to advance, it is essential to consider their broader impacts beyond technical innovation. The large-scale training and deployment of RSFMs raise important environmental and ethical concerns, while also offering the potential for substantial societal benefit [1]. This is particularly relevant given the strategic role of RS in EO, environmental monitoring, disaster response, and national security. A recent effort by Ghamisi et al. [156] provides a timely and comprehensive review on responsible AI for EO. Their work highlights key dimensions including security, privacy, ethical risks, algorithmic bias, and fairness, and proposes achievable pathways to align model development with societal values and public interest. While our survey emphasizes the technical foundations of RSFMs, we refer readers to this work for an in-depth discussion of responsible practices and their relevance to the RS community.
VI-F Summary and Discussion
VII Benchmark Performance
In this section, we compare, analyze, and discuss the existing RSFM studies reviewed in this survey. It is important to note that these studies have distinct research focuses, modalities, learning setups, and target tasks or applications, making it difficult to conduct fair comparisons for all of them using unified benchmarks. Furthermore, as RSFM development is still in its early stages, robust and comprehensive evaluation frameworks across different tasks and modalities are yet to be fully established. Given these limitations, we focus on key dimensions and try to provide an informative analysis and insights into the current progress and future directions.